# UNGER: Generative Recommendation with A Unified Code via Semantic and Collaborative Integration

### v2

@@ Abstract.
With the recent surge in the popularity of generative paradigms, generative recommendation has gradually drawn attention from researchers in the recommendation community. The core component in generative recommendation is the item code, which is generally acquired through the quantization of collaborative or semantic modality representation, and then acts as the candidate item identifier in the auto-regressive generative model for sequence-to-sequence modeling. However, existing methods typically construct independent codes for different modalities, which not only double the computational and storage costs but also impede the effective integration of the complementary strengths of collaborative and semantic knowledge. Considering this limitation, this paper seeks to integrate two different modalities into a unified code, fully unleashing the potential of complementary nature among modalities. Nevertheless, the integration remains challenging: the integrated embedding obtained by the common concatenation method aligns predominantly with semantic representation, overshadowing collaborative information. This imbalance would lead to underutilization of collaborative knowledge, thereby resulting in limited effectiveness.
To address this, we propose a novel method, named UNGER, which integrates semantic and collaborative knowledge into a unified code for generative recommendation. Specifically, we propose to adaptively learn an integrated embedding through the joint optimization of cross-modality knowledge alignment and next item prediction tasks. Subsequently, to mitigate the information loss caused by the quantization process, we further introduce an intra-modality knowledge distillation task, using the integrated embeddings as supervised signals to compensate during the training of the auto-regressive generative model. Extensive experiments on three widely used benchmarks demonstrate the superiority of our approach compared to existing methods, including both generative and traditional paradigms.
Generative Recommendation, Autoregressive Generation, Semantic Tokenization, Collaborative-Semantic Fusion
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Recommendation systems have become widely adopted in modern society, such as in video (Zhao et al., 2019), movie (Diao et al., 2014), and music (Kim et al., 2007) scenarios. Modern recommendation systems typically adopt a retrieve-and-rank strategy. During the retrieval process, the model needs to retrieve the most relevant items from a large candidate pool based on the item embeddings learned through representation learning. However, existing approaches, whether based on dual-tower models (Covington et al., 2016; Wang et al., 2021) or graph models (Wang et al., 2018; Ying et al., 2018), generally rely on dot-product (cosine) similarity with approximate nearest neighbor (ANN) search (Andoni et al., 2018; Arya et al., 1998). Although ANN accelerates top-k recommendation retrieval, its search indices are constructed using tools such as Faiss (Johnson et al., 2019) and SCANN (Guo et al., 2020), which are independent of the optimization process of the recommendation model, thereby limiting the overall effectiveness of recommendation systems (Zhu et al., 2024).
@@
To address this limitation, generative recommendation (Liu et al., 2024c; Wang et al., 2023a, b; Rajput et al., 2023), which adopts code as the core component and frames the recommendation task as autoregressive code sequence generation, has emerged as a promising research direction owing to its potential to enable more efficient decoding without the need for ANN indexes. Unlike traditional methods that match users and items based on embeddings, generative recommendation predicts candidate item code directly, as depicted in Figure 1. Specifically, pre-trained semantic encoders (Ni et al., 2021; Lee and Toutanova, 2018) or collaborative encoders (Zhou et al., 2018; Kang and McAuley, 2018) are first utilized to encode item semantic or ID information into embeddings. Then, embedding quantization (Lee et al., 2022; Qi et al., 2017) is applied to map these embeddings into discrete codes forming a lookup table of item codes. The input of transformer will be converted into a sequence of codes, enabling the model to autoregressively predict the next item’s code step by step.
@@
Existing methods typically construct independent codes for different modalities knowledge, either collaborative or semantic modality. For instance, Recforest (Feng et al., 2022) encodes item collaborative knowledge into a single set of codes, whereas TIGER (Rajput et al., 2023) focuses on encoding item semantic knowledge into a separate code set. Furthermore, EAGER (Wang et al., 2024b) encodes both collaborative and semantic knowledge into two independent sets of codes. These approaches have demonstrated that leveraging both collaborative and semantic information is necessary to enhance the effectiveness and robustness of recommendation models. However, for generative recommenders aiming to integrate both semantic and collaborative knowledge, two separate codes are unacceptable, as the dual-code framework significantly increases storage and inference costs, rendering it impractical for large-scale deployments. Besides, the intrinsic misalignment between semantic and collaborative knowledge limits the potential to fully exploit their complementary strengths with two separate codes (Kim et al., 2024).
@@
However, it is non-trivial to adopt a unified code. The core challenge in adopting a unified code lies in how to get a unified representation which integrates knowledge from both collaborative and semantic modalities. A common practice is to concatenate semantic and collaborative features to create a unified representation. However, this approach suffers from a phenomenon termed the semantic dominance issue. As depicted in Figure 3, the final representation aligns more closely with the semantic side, leading to underutilization of the collaborative side. This imbalance not only limits the potential of integrating collaborative knowledge but, in extreme cases, can even degrade performance to levels below those achieved with semantic knowledge alone. The root cause of this issue lies in the inability of concatenation to account for the inherent differences between the two different modalities, such as variations in signal strength and representational misalignment (Wang et al., 2024a). Consequently, semantic representations dominate the final output, undermining the expected advantages of unified representations.
@@
Firstly, a unified code that encodes both collaborative and semantic knowledge is crucial for practical deployment. Due to the inherent modality differences (Kim et al., 2024; Wang et al., 2024a), adopting a unified code poses challenges. Therefore, we propose a two-stage knowledge integration framework. Specifically, in the first stage, it uses Llama2-7b (Touvron et al., 2023) to extract semantic knowledge and DIN (Zhou et al., 2018) to extract collaborative knowledge, then adaptively learn a unified representation which integrates both knowledge to construct discrete item codes, referred to as Unicodes, which serve as item identifiers for the second stage. In the second stage, our architecture consists of an encoder that models the user’s interaction history and a decoder that predicts the unicode of the target item.
Secondly, to resolve the semantic domination issue, we propose a learnable modality adaption layer and a cross-modality knowledge alignment task, which help to adaptively learn an integrated embedding through the joint optimization with next item prediction task. By providing fine-grained contrastive supervision signals, it encourages the representation to leverage both collaborative and semantic knowledge more balanced, avoiding the overemphasis on semantic modality.
Thirdly, ensuring comprehensive and sufficient learning is essential for fully exploiting both knowledge. To compensate for the information loss caused by the quantization process, we introduce an intra-modality knowledge distillation task with a special token in the second stage. Inspired by contrastive learning paradigms and the success of the [CLS] token in BERT (Devlin, 2018) for classification tasks, we design a special token to summarize the sequence’s knowledge. The embedding of this token is then used in a global contrastive task against the target item integrated embedding, enabling the model to distill the integrated knowledge.
@@ 2. RELATED WORK
For LLMs based methods (Zheng et al., 2024; Lin et al., 2024; Liu et al., 2024b; Kim et al., 2024; Ji et al., 2024), the focus is on designing refined prompts and fine-tuning tasks that help language models better understand recommendation tasks. LC-Rec (Zheng et al., 2024) introduces a learning-based vector quantization approach for semantically meaningful item indexing and fine-tuning tasks that align collaborative signals with LLM representations, achieving superior performance in diverse scenarios. CCF-LLM (Liu et al., 2024b) transforms user-item interactions into hybrid prompts encoding both semantic knowledge and collaborative signals, utilizing an attentive cross-modal fusion strategy to integrate embeddings from different modalities. SC-Rec (Kim et al., 2024) utilizes multiple item indices and prompt templates, alongside a self-consistent re-ranking mechanism, to more effectively merge collaborative and semantic knowledge.
For methods that train from scratch (Zeng et al., 2024; Wang et al., 2024a; Liu et al., 2024a; Feng et al., 2022; Rajput et al., 2023; Wang et al., 2024b; Zhu et al., 2018, 2019), the primary focus is on converting the raw sequence recommendation task into an autoregressive generation task. Tree-based methods (Zhu et al., 2018; Feng et al., 2022; Zhu et al., 2019), such as RecForest (Feng et al., 2022), have shown promising performance by constructing multiple trees and integrating transformer-based structures for routing. Additionally, TIGER (Rajput et al., 2023) introduced the concept of semantic IDs, representing each item as a set of tokens derived from its side information, and predicting next item tokens in a sequence-to-sequence manner. EAGER (Wang et al., 2024b) employs a dual-stream generative framework to parallely utilize semantic and behavioral information with two separate codes, generating recommended items from each respective pipeline, and ultimately selecting the top-k items based on confidence scores.
@@ 3.3. Stage I: Item Unicodes Generation
Modality adaption layer. To bridge the modality gap between two modalities, we propose using a learnable modality adaptation layer with AdaLN that maps the semantic embeddings $E_{S}$ into the same embedding space as collaborative embeddings. AdaLN (Peebles and Xie, 2023) introduces learnable parameters to adaptively adjust the normalization process according to the input, effectively capturing the distinctive characteristics of the two different modalities. The mapping process is defined as follows:
@@
Cross-modality knowledge alignment task. To effectively integrate semantic and collaborative knowledge, we introduce a cross-modality knowledge alignment task designed to align embeddings from the collaborative modality $E_{C}$ and semantic modality $E_{T}$. For a given item $i\in I$, our objective is to bring its collaborative embedding $E_{C_{i}}$ closer to its semantic embedding $E_{T_{i}}$ for item $i$ while distancing $E_{C_{i}}$ from the semantic embedding $E_{T_{j}}$ of a different item $j\in I$, where $j$ is an in-batch negative sample selected from the same batch as item $i$. By structuring the alignment in this way, the learned embedding for item $i$ is able to encapsulate knowledge from both modalities. To achieve this goal, we adopt the Info-NCE loss (Chen et al., 2020). The loss function for our cross-modality knowledge alignment task is defined as follows:
@@
Next item prediction task. In addition to the alignment loss described above, we adopt next item prediction task to enhance the model’s ability to capture the sequence dependency, thereby supporting the learning of robust collaborative representations. This task takes the user’s historical interaction sequence $[x_{1},x_{2},\dots,x_{t-1}]$ as input, learns a representation of the user’s preferences, and computes matching scores by measuring the similarity between the learned preferences and candidate items. These scores are subsequently normalized to derive the probability distribution over the target item. The goal is to maximize the probability assigned to the actual target item $x_{t}$, which can be formally represented as follows:
@@
Joint optimization. In the first stage, we jointly optimize the two losses to train the modality adaption layer and sequential recommendation model. The total loss function for the first stage is defined as follows:

### v3

@@ Abstract.
With the rise of generative paradigms, generative recommendation has garnered increasing attention. The core component is the item code, generally derived by quantizing collaborative or semantic representations to serve as candidate items identifiers in the context. However, existing methods typically construct separate codes for each modality, leading to higher computational and storage costs and hindering the integration of their complementary strengths. Considering this limitation, we seek to integrate two different modalities into a unified code, fully unleashing the potential of complementary nature among modalities. Nevertheless, the integration remains challenging: the integrated embedding obtained by the common concatenation method would lead to underutilization of collaborative knowledge, thereby resulting in limited effectiveness.
To address this, we propose a novel method, named UNGER, which integrates semantic and collaborative knowledge into a unified code for generative recommendation. Specifically, we propose to adaptively learn an integrated embedding through the joint optimization of cross-modality knowledge alignment and next item prediction tasks. Subsequently, to mitigate the information loss caused by the quantization process, we introduce an intra-modality knowledge distillation task, using the integrated embeddings as supervised signals to compensate. Extensive experiments on three widely used benchmarks demonstrate the superiority of our approach compared to existing methods.
Generative Recommendation, Autoregressive Generation, Semantic Tokenization, Collaborative-Semantic Integration
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/04ccs: Information systems Recommender systems
@@ 1. Introduction
To mitigate the issue of information overload, recommendation systems have become widely adopted in modern society, such as in video (Zhao et al., 2019), movie (Diao et al., 2014), and music (Kim et al., 2007) scenarios. Recommendation systems mine users’ interests and proactively recommend items that they may be interested in. Modern recommendation systems typically adopt a retrieve-and-rank strategy. During the retrieval process, the model needs to retrieve the most relevant items from a large candidate pool based on the item embeddings learned through representation learning. However, existing approaches, whether based on dual-tower models (Covington et al., 2016; Wang et al., 2021) or graph models (Wang et al., 2018; Ying et al., 2018), generally rely on dot-product (cosine) similarity with approximate nearest neighbor (ANN) search (Andoni et al., 2018; Arya et al., 1998). Although ANN accelerates top-k recommendation retrieval, its search indices are constructed using tools such as Faiss (Johnson et al., 2019) and SCANN (Guo et al., 2020), which are independent of the optimization process of the recommendation model, thereby limiting the overall effectiveness of recommendation systems (Zhu et al., 2024).
@@
To address this limitation, generative recommendation (Liu et al., 2024c; Wang et al., 2023a, b; Rajput et al., 2023), which adopts code as the core component and frames the recommendation task as autoregressive code sequence generation, has emerged as a promising research direction owing to its potential to enable more efficient decoding without the need for ANN indexes. Unlike traditional methods that match users and items based on embeddings, generative recommendation predicts candidate item code directly, as depicted in Figure 1. Specifically, pre-trained semantic encoders (Ni et al., 2021; Lee and Toutanova, 2018) or collaborative encoders (Zhou et al., 2018; Kang and McAuley, 2018) are first utilized to encode item semantic or ID information such as shown in Figure 2 into embeddings. Then, embedding quantization (Lee et al., 2022; Qi et al., 2017; Li et al., 2024) is applied to map these embeddings into discrete codes forming a lookup table of item codes. The input of transformer will be converted into a sequence of codes, enabling the model to autoregressively predict the next item’s code step by step.
Figure 2. An example of item semantic knowledge.
@@
Existing methods typically construct independent codes for different modalities knowledge, either collaborative or semantic modality. For instance, Recforest (Feng et al., 2022) encodes item collaborative knowledge into a single set of codes, whereas TIGER (Rajput et al., 2023) focuses on encoding item semantic knowledge into a separate code set. Furthermore, EAGER (Wang et al., 2024c) encodes both collaborative and semantic knowledge into two independent sets of codes. These approaches have demonstrated that leveraging both collaborative and semantic information is necessary to enhance the effectiveness and robustness of recommendation models. However, for generative recommenders aiming to integrate both semantic and collaborative knowledge, two separate codes are unacceptable, as the dual-code framework significantly increases storage and inference costs, rendering it impractical for large-scale deployments. Besides, the intrinsic misalignment between semantic and collaborative knowledge limits the potential to fully exploit their complementary strengths with two separate codes (Kim et al., 2024).
@@
However, it is non-trivial to adopt a unified code. The core challenge in adopting a unified code lies in how to get a unified representation which integrates knowledge from both collaborative and semantic modalities. A common practice is to concatenate semantic and collaborative features to create a unified representation. However, this approach suffers from a phenomenon termed the semantic dominance issue. As depicted in Figure 4, the final representation aligns more closely with the semantic side, leading to underutilization of the collaborative side. This imbalance not only limits the potential of integrating collaborative knowledge but, in extreme cases, can even degrade performance to levels below those achieved with semantic knowledge alone. The root cause of this issue lies in the inability of concatenation to account for the inherent differences between the two different modalities, such as variations in signal strength and representational misalignment (Wang et al., 2024b). Consequently, semantic representations dominate the final output, undermining the expected advantages of unified representations.
@@
Firstly, a unified code that encodes both collaborative and semantic knowledge is crucial for practical deployment. In real-world recommendation systems, storing and retrieving items efficiently requires a compact and discrete representation that integrates all essential information related to items. However, due to the inherent differences between collaborative and semantic modalities (Kim et al., 2024; Wang et al., 2024b), directly combining them often leads to suboptimal representations. Semantic embeddings derived from large language models typically contain rich contextual signals, while collaborative embeddings reflect users’ actual preferences through interaction histories. To this end, we propose a two-stage knowledge integration framework. In the first stage, we separately extract modality-specific knowledge using two pretrained models. Specifically, we adopt Llama2-7b (Touvron et al., 2023) to obtain semantic representations based on textual descriptions and DIN (Zhou et al., 2018) to model collaborative behavior patterns. These modality-specific embeddings are then fused through a modality-adaptive fusion module, which is designed to learn an integrated representation that balances and aligns the semantic and collaborative knowledge. The integrated embeddings are subsequently quantized into discrete item codes, referred to as Unicodes, which act as universal identifiers for each item in our model. In the second stage, we utilize the Unicodes within a sequence modeling architecture for recommendation. An encoder is employed to capture the user’s interaction history, while a decoder predicts the Unicode corresponding to the next item. This reformulation allows the recommendation task to be viewed as a discrete sequence generation task.
Secondly, to resolve the semantic domination issue, we introduce two components: a learnable modality adaptation layer and a cross-modality knowledge alignment task. The adaptation layer adjusts the relative emphasis of each modality during the fusion process, allowing the model to modulate their contributions dynamically. Meanwhile, the knowledge alignment task provides explicit supervision by aligning the distributions of semantic and collaborative embeddings through a contrastive learning objective. By encouraging the representations of the same item across modalities to be close in the latent space, the model is guided to capture shared semantics while preserving modality-specific features. This alignment is jointly optimized with the next-item prediction objective, allowing the model to benefit from both modality alignment and end-task supervision. This joint training encourages a more balanced representation that leverages complementary knowledge and reduces the risk of over-reliance on semantic features.
Thirdly, ensuring comprehensive and sufficient learning is essential for fully exploiting both knowledge. The quantization process, although necessary for obtaining discrete Unicodes, inevitably introduces approximation errors that may lead to partial information loss. Since the quantized codes are derived by approximating the original embeddings, relying solely on them may result in insufficient representation of the items. To compensate for this, we introduce an auxiliary learning objective based on intra-modality knowledge distillation, where the original embeddings are leveraged to provide more complete and informative guidance. Motivated by the effectiveness of the [CLS] token in BERT (Devlin, 2018), we incorporate a special token into the input sequence during the second stage. This token serves to summarize the sequence’s informative content. We then conduct a global contrastive learning task between the embedding of this token and the integrated embedding of the target item. This distillation task enables the model to capture and retain high-level modality-specific knowledge that might be attenuated during quantization. By maximizing agreement between the sequence summary and the target item’s unified representation, the model learns to distill and refine knowledge more effectively.
@@ 2. RELATED WORK
In this section, we review the relevant literature and highlight the distinctions between our approach and the existing studies.
2.1. Pre-training in Recommender Systems
Similar to the field of Natural Language Processing (NLP), the advancement of recommendation models also encounters challenges due to the data sparsity (Pan et al., 2019). Traditional recommendation approaches depend heavily on extensive user–item interaction records to model both user preferences and item attributes, as well as the latent relationships between users and items, but when new users or new items come in with few or no interaction histories, these methods are struggle to model their latent representations accurately, resulting in poor recommendation quality and giving rise to the cold-start problem in recommender systems (Lika et al., 2014; Lam et al., 2008; Li et al., 2025), which significantly restricts the scalability and efficacy of classical recommendation paradigms.
To address this challenge, researchers have increasingly turned to pre-training strategies inspired by breakthroughs in NLP. Early methods primarily focus on learning universal user representations from large-scale interaction data, often using self-supervised or multi-task learning objectives based on ID-level signals. For example, PeterRec (Yuan et al., 2020) enables knowledge transfer across multiple downstream tasks through parameter-efficient fine-tuning, while Conure (Yuan et al., 2021) introduces a lifelong learning framework that incrementally updates user embeddings without forgetting previous knowledge. CLUE (Cheng et al., 2021) further improves representation quality by adopting contrastive learning at the sequence level, bridging the gap between item-level objectives and user-level semantics. However, such ID-based approaches struggle with transferability when applied to new domains or scenarios, as they rely heavily on discrete user and item identifiers that lack generalization capability. To address this, several work incorporate rich side information such as item descriptions and user reviews to enhance semantic understanding. ZESRec (Ding et al., 2021) replaces item IDs with textual metadata to enable zero-shot generalization across entirely disjoint domains, while UniSRec (Hou et al., 2022) and MISSRec (Wang et al., 2023c) leverage multimodal features including text and images to build transferable sequence and item representations that better reflect user preferences. Building upon these advances, P5 (Geng et al., 2022) unifies various recommendation tasks into a text-to-text framework, transforming all inputs and outputs into natural language and enabling a single language model to handle diverse recommendation objectives with minimal adaptation. Nevertheless, these approaches still rely on small-scale pretrained language models, which constrains their capacity to capture universal world knowledge and provide general representation.
2.2. Sequential Recommendation
@@
For LLMs based methods (Zheng et al., 2024; Lin et al., 2024; Liu et al., 2024b; Shen et al., 2024; Kim et al., 2024; Ji et al., 2024), the focus is on designing refined prompts and fine-tuning tasks that help language models better understand recommendation tasks. LC-Rec (Zheng et al., 2024) introduces a learning-based vector quantization approach for semantically meaningful item indexing and fine-tuning tasks that align collaborative signals with LLM representations, achieving superior performance in diverse scenarios. CCF-LLM (Liu et al., 2024b) transforms user-item interactions into hybrid prompts encoding both semantic knowledge and collaborative signals, utilizing an attentive cross-modal fusion strategy to integrate embeddings from different modalities. SC-Rec (Kim et al., 2024) utilizes multiple item indices and prompt templates, alongside a self-consistent re-ranking mechanism, to more effectively merge collaborative and semantic knowledge.
For methods that train from scratch (Zeng et al., 2024; Wang et al., 2024b; Liu et al., 2024a; Feng et al., 2022; Rajput et al., 2023; Wang et al., 2024c; Zhu et al., 2018, 2019; Wang et al., 2024a), the primary focus is on converting the raw sequence recommendation task into an autoregressive generation task. Tree-based methods (Zhu et al., 2018; Feng et al., 2022; Zhu et al., 2019), such as RecForest (Feng et al., 2022), have shown promising performance by constructing multiple trees and integrating transformer-based structures for routing. Additionally, TIGER (Rajput et al., 2023) introduced the concept of semantic IDs, representing each item as a set of tokens derived from its side information, and predicting next item tokens in a sequence-to-sequence manner. EAGER (Wang et al., 2024c) employs a dual-stream generative framework to parallely utilize semantic and behavioral information with two separate codes, generating recommended items from each respective pipeline, and ultimately selecting the top-k items based on confidence scores.
@@ 3.3. Stage I: Item Unicodes Generation
3.3.1. Modality adaption layer
To bridge the modality gap between two modalities, we propose using a learnable modality adaptation layer with AdaLN that maps the semantic embeddings $E_{S}$ into the same embedding space as collaborative embeddings. Unlike standard Layer Normalization, AdaLN (Peebles and Xie, 2023) introduces learnable affine parameters that are conditioned on the input itself, enabling the model to dynamically adjust the normalization process based on the characteristics of each modality. This design choice is particularly well-suited for our setting, where semantic and collaborative embeddings inherently capture distinct patterns (e.g., content-driven semantics vs. behavior-driven preferences). By leveraging AdaLN, the adaptation layer is better equipped to preserve and align these modality-specific signals. The mapping process is defined as follows:
@@ 3.3.1. Modality adaption layer
3.3.2. Cross-modality knowledge alignment task
To effectively integrate semantic and collaborative knowledge, we introduce a cross-modality knowledge alignment task designed to align embeddings from the collaborative modality $E_{C}$ and semantic modality $E_{T}$. For a given item $i\in I$, our objective is to bring its collaborative embedding $E_{C_{i}}$ closer to its semantic embedding $E_{T_{i}}$ for item $i$ while distancing $E_{C_{i}}$ from the semantic embedding $E_{T_{j}}$ of a different item $j\in I$, where $j$ is an in-batch negative sample selected from the same batch as item $i$. By structuring the alignment in this way, the learned embedding for item $i$ is able to encapsulate knowledge from both modalities. To achieve this goal, we adopt the Info-NCE loss (Chen et al., 2020). The loss function for our cross-modality knowledge alignment task is defined as follows:
@@ 3.3.2. Cross-modality knowledge alignment task
3.3.3. Next item prediction task
In addition to the alignment loss described above, we adopt next item prediction task to enhance the model’s ability to capture the sequence dependency, thereby supporting the learning of robust collaborative representations. This task takes the user’s historical interaction sequence $x_{1},x_{2},\dots,x_{t-1}]$ as input, learns a representation of the user’s preferences, and computes matching scores by measuring the similarity between the learned preferences and candidate items. These scores are subsequently normalized to derive the probability distribution over the target item. The goal is to maximize the probability assigned to the actual target item $x_{t}$, which can be formally represented as follows:
@@ 3.3.3. Next item prediction task
3.3.4. Joint optimization
In the first stage, we jointly optimize the two losses to train the modality adaption layer and sequential recommendation model. The total loss function for the first stage is defined as follows: